Bright Star School purchases an activity box from a local supplier in quantity (D) of 18000 boxes per annum. Annual carrying cost (H) is 5 % of the price of box and ordering cost (S) is AED 200 per order. The school is open 250 days a year. The price of each box is AED 220. Apply the Economic Order quantity model of Inventory management and answer the following:   

  1. Determine the optimal number of the items to order in one lot i.e. Economic Order Quantity (EOQ).                             
  2. Determine number of orders per year?
  3. Expected Time between orders?
  4. Total annual cost (Ordering cost + Holding cost)?
  5. The supplier is ready to offer a discount of 10 % if a bulk order of 2000 items is given. Analyze whether this discount offer will be beneficial for the company or not.
Answer #1

EOQ = [ 2 x D xS /Px H]1/2

= [ 2 x 18000 x 200 /220x0.05]1/2

= 809

Number of orders = annual demand / EOQ = 18000/809 = 22.24

Time beteeen orders = Number of days / number of orders = 250/22.24 =11.241

Total cost = ordering cost + holding cost

= 22.24 x200 + 809/2 x11 = 8898

Note : As per Chegg's policy, I am authorised to answer the first four parts of a problem only. Inconvenience is regretted.